OpenStudy (texmechanica):

Consider the system of equations: 3x+4y=67 and 5x-y=35 Edmundo used the equations in the system to find another equation whose graph would pass through the solution of the system of equations. What number completes this equation? 8x+3y=?

OpenStudy (faiqraees):

add both equations 3x+4y=67 5x-y=35 ---------- 8x+3y= 102